Water is added to the solution such that the mole fraction of water in the solution becomes 0.9. The boiling point of the solution is 
a.375.5K        b.376.2K     c.380.4K

Grade:

1 Answers

Arun
4 years ago
Dear student
 
X(H2O) = 0.9
X (C2H5OH) = 0.1
 
delta Tb = Kb * m
0.52 * 0.1 * 1000/ (0.9 * 18) = 3.21 K
 
hence boiling point Tb = 373 + 3.21 = 376.21K